OpenStudy (anonymous):

What is the greatest common factor of 19x^7 and 3x^5? A. x5 B. x7 C. 3x5 D. 3x7

OpenStudy (anonymous):

Are you here to learn or just trying to get someone to do your homework for you?

OpenStudy (anonymous):

Uhh im trying to get help with my problems ? Are you on here to help or criticize people ?

OpenStudy (uri):

Well i have seen you asking these same type of question 3 times.@Babez_1999 What do you think the answer will be?

OpenStudy (anonymous):

I ask the question because this one is exactly the same as your last, yet you have shown no attempt at answering it. Break the terms down into the prime factors.

OpenStudy (anonymous):

@uri it's the plan that im working on in the work book , & @gypsy1274 i don't need you help (: Buuhh Bye and i wasn't taught this is school this is my frist time ever doing this so yeah if i don't understand it the first 2 times im gonna keep asking until i can figure it out .

OpenStudy (anonymous):

I can help you, I just wanted to be clear about what you are asking for.

OpenStudy (anonymous):

like i said i don't need your help @gypsy1274

OpenStudy (uri):

@Babez_1999 Be nice.

OpenStudy (mathstudent55):

@Babez_1999 Do the same we did before. Try to find the GCF of the numbers. Then try to find the GCF of the varuiables. When two numbers have no common factor, you can say that their GCF is 1. Then multiply the GCF of the numbers by the GCF of the variables to get your answer.

OpenStudy (anonymous):

i just don't get when i have to do prime factorization 19 = 8 and 3=2 ? that's what i got but i don't know where to go from there

OpenStudy (mathstudent55):

19 is a prime number. It can't be factored. 3 is also a prime number. It can't be factored. Every time you need to find the GCF of two numbers, you can do prime factorization of both numbers. That method will always give you the GCF. Sometimes, the numbers are so simple, that you can just look at them and find the GCF easily without prime factorization. For example, the GCF of 6 and 10 is 2. You can just look at 6 and 10 and conclude that. Using prime factorization, you'd do: 6 = 2 * 3 10 = 2 * 5 Which factors do they have in common? Only 2, so 2 is the GCF of 6 and 10.

OpenStudy (mathstudent55):

Back to this question. The prime factorizations of 19 and 3 are: 19 = 19 3 = 3 The GCF of 19 and 3 is 1. The GCF of x^7 and x^5 is x^5 The GCF of 19x^7 and 3x^5 is 1 * x^5 = x^5